when a force is applied at one end of an elastic wire it produce a strain e in the wire . if the y is the young modulus of the material of the wire what is the amount of energy stored per unit volume of the wire
1
Expert's answer
2018-03-20T10:45:08-0400
Energy stored per unit volume,
E =1/2 σ ε
We know that,
σ=εY
So, we get
E =1/2 Y〖 ε〗^2
Answer: E =1/2 Y〖 ε〗^2